If the coefficient of kinetic friction between a 33.0 kg crate and the floor is 0.360, what horizontal force is required to move the crate at a steady speed across the floor?
Answer #1

from the newton second law

net force = ma

here crate is moving with a constant speed so acceleration is zero

net force = 0

F-fk = 0

F = fk = muek*n

n = normal force = mg

F = 0.36*33*9.8 = 116.4 N